OpenStudy (anonymous):

x-y=-2 2x-3y=2 solve my substituting.

OpenStudy (jiteshmeghwal9):

First find x from one of the equation I gt x from Ist equation x=-2+y Put this value in iind equation i.e; 2(-2+y)-3y=2 Solve for y & then put the value of y in the equation x=-2+y to gt x :)

OpenStudy (anonymous):

so i got -4+2y-3y=2 thats where i get stuck :(

OpenStudy (jiteshmeghwal9):

-4-y=2 y=-4-2=-6

OpenStudy (jiteshmeghwal9):

now put the value of y in the equation x=-2+y to gt x :)

OpenStudy (anonymous):

how did the 6 become negative?

OpenStudy (jiteshmeghwal9):

-4-2=-(4+2)=-6, gt it ?

OpenStudy (anonymous):

oh okay. thank you so much.

OpenStudy (hba):

Do you want to check whether your answers are right or wrong ?

OpenStudy (jiteshmeghwal9):

Easy put the values in the equation & conform ur answer :)